A 10 kg rock is thrown up. It is moving at 9 m/s. What is its potential energy at the top? (Hint: conservation of energy)
1
Expert's answer
2018-05-04T12:51:08-0400
The law of conservation of energy states
KE+PE=const
So
PE_final=KE_initial
PE_final=(mv_i^2)/2=(10×9^2)/2=405 J